Question:

Find the equation of the tangent plane and normal to the surface

2π‘₯𝑧2 βˆ’ 3π‘₯𝑦 βˆ’ 4π‘₯ = 7 at point

( 1, βˆ’1,2 )



1
Expert's answer
2021-08-12T14:09:43-0400

Solution;

Now, as gradient of 𝑓 is normal at a point to the level surface, we have;

"\nabla f=\tilde{i}\frac{\delta f}{\delta x}+\tilde{j}\frac{\delta f}{\delta y}+\tilde{k}\frac{\delta f}{\delta z}"

Since , "f=2xz^2-3xy-4x-7"

We have;

"\frac{\delta f}{\delta x}=2z^2-3y-4"

"\frac{\delta f}{\delta y}=-3x"

"\frac{\delta f}{\delta z}=4xz"

Therefore;

"\nabla f=\tilde{i}(2z^2-3y-4)+\tilde{j}(-3x)+\tilde{k}(4xz)"

At point (1,-1,2),we have the gradient as;

"\nabla f=\tilde{i}(2(2^2)-3(-1)-4)+\tilde{j}(-3×1)+\tilde {k}(4×1×2)"

"\nabla f=7\tilde i-3\tilde j +8\tilde k" Is the tangent.

If the vector position at point (1,-1,2) is "\vec r" then we have;

"\vec r=\tilde i-\tilde j+2\tilde k"

Let Q be another current point on the level surface through P(1,-1,2) with position vector "\vec R" .

Then ;

"\vec {PQ}" ="\vec R-\vec r" ="(X-1)\tilde i +(Y+1)\tilde j+(Z-2)\tilde k"

Since the tangent "\nabla f" and "\vec {PQ}" are perpendicular vectors ,their dot product is 0,we have

"[(X-1)\tilde i+(Y+1)\tilde I+(Z-2)\tilde k].[7\tilde i -3\tilde j+8\tilde k]=0"

Hence;

7(X-1)-3(Y+1)+8(Z-2)=0

7X-3Y+8Z=26

If the position vector of the current point Q on the normal is "\vec R" and (X,Y,Z) are its coordinates,then the equation of the normal at point P(x,y,z) to a level surface is given by

"\frac{X-x}{\frac{\delta f}{\delta x}}=\frac{Y-y}{\frac{\delta f}{\delta y}}=\frac{Z-z}{\frac{\delta f}{\delta z}}"

Here,point (x,y,z) is (1,-1,2) and "\frac{\delta f}{\delta x}=7" ,"\frac{\delta f}{\delta y}=-3" and "\frac{\delta f}{\delta z}=8"

Therefore,the equation of the normal at the point (1,-1,2) to a given level surface is

"\frac{X-1}{7}=\frac{Y+1}{-3}=\frac{Z-2}{8}"